# Homological stratification and descent

### 04 Mar 2026 v2

@@ 11. Equivariant applications
induced by base-change $S_{G}^{0}\to R$ and the $H$-geometric fixed point functors for $\operatorname{Mod}_{G}(R)$ and $\operatorname{Sp}_{G}$. This diagram shows that the images of the maps $\varphi^{h}_{H}$ are disjoint in $\operatorname{Spc}^{h}(\operatorname{Mod}_{G}(R)^{c})$ for non-conjugate subgroups, since this is true for the $R=S^{0}_{G}$ case just mentioned. As a consequence of 11.5 and 7.9, these images also cover the homological spectrum, which gives the desired decomposition of the homological spectrum. The analogous statement for the tt-spectrum is proved similarly. ∎
@@ 12. Open questions
We have defined h-stratification for any rigidly-compactly generated tt-category (3.2), but we only defined tt-stratification for those $\mathscr{T}$ whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian (8.5). This is because the latter topological condition is needed for the construction of the Balmer–Favi support. However, W. Sanders [SAN17] has introduced a generalization of the Balmer–Favi support which does not require the spectrum to be weakly noetherian. The caveat is that the Balmer–Favi–Sanders support of an object (or localizing ideal) is always closed with respect to the so-called localizing topology on $\operatorname{Spc}(\mathscr{T}^{c})$; see [SAN17, Theorem 4.2]. [^4] A spectral space is weakly noetherian if and only if its localizing topology is discrete. See [ZOU23, Section 2] for further information about the localizing topology. Using this theory of support, one can extend the definition of tt-stratification to arbitrary rigidly-compactly generated tt-categories as in [ZOU23, Definition 8.1]. However, it turns out that if $\mathscr{T}$ is tt-stratified in this a priori more general sense, then the spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is necessarily weakly noetherian; see [ZOU23, Theorem 8.13]. Thus, using the Balmer–Favi–Sanders support does not provide a more general notion of tt-stratification. With this and 8.6 in mind, it is natural to ask:
@@
Suppose $\mathscr{T}$ is a rigidly-compactly generated tt-category. If $\mathscr{T}$ is h-stratified and satisfies the Nerves of Steel conjecture, does it follow that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian?
@@
Example 12.12.
By 1.15 this holds whenever $\mathscr{T}$ has enough tt-fields. For example, this holds in the derived category $\operatorname{D}(R)$ of a commutative ring.

### 18 Mar 2026 v3

@@ 11. Equivariant applications
induced by base-change $S_{G}^{0}\to R$ and the $H$-geometric fixed point functors for $\operatorname{Mod}_{G}(R)$ and $\operatorname{Sp}_{G}$. This diagram shows that the images of the maps $\varphi^{h}_{H}$ are disjoint in $\operatorname{Spc}^{h}(\operatorname{Mod}_{G}(R)^{c})$ for non-conjugate subgroups, since this is true for the $R=S^{0}_{G}$ case just mentioned. As a consequence of 11.5 and 7.9, these images also cover the homological spectrum, which gives the desired decomposition of the homological spectrum. The analogous statement for the $t$-spectrum is proved similarly. ∎
@@ 12. Open questions
We have defined h-stratification for any rigidly-compactly generated $t$-category (3.2), but we only defined $t$-stratification for those $\mathscr{T}$ whose spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian (8.5). This is because the latter topological condition is needed for the construction of the Balmer–Favi support. However, W. Sanders [SAN17] has introduced a generalization of the Balmer–Favi support which does not require the spectrum to be weakly noetherian. The caveat is that the Balmer–Favi–Sanders support of an object (or localizing ideal) is always closed with respect to the so-called localizing topology on $\operatorname{Spc}(\mathscr{T}^{c})$; see [SAN17, Theorem 4.2]. [^4] A spectral space is weakly noetherian if and only if its localizing topology is discrete. See [ZOU23, Section 2] for further information about the localizing topology. Using this theory of support, one can extend the definition of $t$-stratification to arbitrary rigidly-compactly generated $t$-categories as in [ZOU23, Definition 8.1]. However, it turns out that if $\mathscr{T}$ is $t$-stratified in this a priori more general sense, then the spectrum $\operatorname{Spc}(\mathscr{T}^{c})$ is necessarily weakly noetherian; see [ZOU23, Theorem 8.13]. Thus, using the Balmer–Favi–Sanders support does not provide a more general notion of $t$-stratification. With this and 8.6 in mind, it is natural to ask:
@@
Suppose $\mathscr{T}$ is a rigidly-compactly generated $t$-category. If $\mathscr{T}$ is h-stratified and satisfies the Nerves of Steel conjecture, does it follow that $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian?
@@
Example $12.$
By 1.15 this holds whenever $\mathscr{T}$ has enough $t$-fields. For example, this holds in the derived category $\operatorname{D}(R)$ of a commutative ring.